A 15.0 kg object is moving with a velocity of +7.50 m/s. A force of -10.0 N acts on the object and its velocity becomes 3.20 m/s. What is the displacement of the object while the force acts?
1
Expert's answer
2013-01-14T11:08:41-0500
The work done by a constant force of magnitude F ona point that moves a distance d in the direction of the forceis the product,
A = F*d
The law of conservationof energy:
T1 + A = T2
T - kinetic energy
T = mv^2/2
A = T2 - T1 = m/2 (v2^2- v1^2)
v1= 7.5 m/s, v2 = 3.2 m/s
d= A/F = m/(2F) (v2^2 - v1^2) =15/(2*10) * (7.5^2 - 3.2^2) = 34.5 m